A $1,000 face value, 240-day bond is quoted at a bank discount yield of 3.3 percent. What is the current bond price? Multiple Choice $957.60 $960.09 $978.00 $982.02 $988.73
Answer #1

Bank Discount Yield (BDY) = (Face Value-Issue Price)/Face Value *360/t

.033 = (1000-P)/1000 * 360/240

(1000-P)/1000 = .033*240/360

= 0.022

1000-P = 0.022*1000

= 22

P = 1000-22

=  $978.00